When 6.13 g iron filings (Fe) are treated with excess hydrochloric acid (aq. HCl(g)), a homogeneous solution of iron(II) chloride (FeCl2(aq)) is formed. Alkalization of this colorless solution results in the formation of iron(II) hydroxide [Fe(OH)2], a white solid that rapidly turns green in the presence of oxygen from the air. When kept under anaerobic conditions, however, iron(II) hydroxide decomposes to form magnetite (Fe3O4), a magnetic solid and binary compound of iron and oxygen. In the space provided below, calculate the mass of oxygen expected in a sample of magnetite prepared as described above.
1
Expert's answer
2020-11-22T11:15:22-0500

Fe + 2HCl = FeCl2 + H2

FeCl2 + 2NaOH = Fe(OH)2 + 2NaCl

4Fe(OH)2 + O2 + 2H2O = 4Fe(OH)3

Fe(OH)2 = FeO + H2O

2Fe(OH)3= Fe2O3 + 3H2O

FeO+Fe2O3=Fe3O4.

this reaction can be written in the form 3Fe = Fe3O4